# Ab Initio Theory of Large Phonon Magnetic Moments Induced by Electron-Phonon Coupling in Magnetic Materials

### v3

@@ Abstract
Through first-principles calculations guided by the inertially decoupled model, we identified five hexagonal ($P6/mmm$) magnets exhibiting intrinsic phonon Chern states, namely, $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ [59], $\mathrm{EuGa_{2}}$, $\mathrm{GdGa_{2}}$ [60], and $\mathrm{CrB_{2}}$ [61, 62]. Remarkably, they all have large observable topological phonon gaps at the $\mathrm{K}$ point defined as $\Delta_{\text{ph}}=\Delta_{\text{z}}-\gamma$, where $\Delta_{\text{z}}=2|E_{\text{Z}}|$ is the phonon Zeeman splitting and $\gamma$ is the phonon linewidth resulting from EPC [63, 64], as summarized in Table I. Note that the Zeeman splitting should be larger than the phonon linewidth to produce an observable phonon gap.
@@
Here, we take $\mathrm{EuSi_{2}}$ as a representative example (see SM [44] for details on all candidate materials). The lattice structures of $\mathrm{EuSi_{2}}$ are illustrated in Figs. 3(a, b). Nonmagnetic $\mathrm{Si}$ atoms form a honeycomb-lattice layer, stacked alternately with a ferromagnetic $\mathrm{Eu}$ triangular-lattice layer, along the [001] direction. The atomic masses of $\mathrm{Si}$ and $\mathrm{Eu}$ are approximately $28\,u$ and $152\,u$ ($u$ is atomic mass unit), respectively, with each $\mathrm{Eu}$ atom carrying an out-of-plane magnetic moment of $7\,\mu_{B}$. As a result, $\mathrm{EuSi_{2}}$ can be aptly described by our proposed inertially decoupled model.
@@
Our first-principles calculations reveal that the EPC induces a pronounced Zeeman splitting of chiral phonons in the high-frequency phonon dispersion of ferromagnetic $\mathrm{EuSi_{2}}$ [see Figs. 3(c, d)], opening a sizable full topological gap ($\approx 0.35$ THz) throughout the Brillouin zone. Given that the Chern number remains $|C|=1$ for all 2D $k_{z}$ slices, $\mathrm{EuSi_{2}}$ realizes a 3D layered phonon Chern state. This is further supported by the presence of boundary modes in the phonon dispersions of $\mathrm{EuSi_{2}}$ under open-boundary conditions, as illustrated in Fig. 3(e). To achieve a 2D Chern state, we construct a dynamically-stable sandwiched slab featuring a $\mathrm{Si}$ sublayer between two $\mathrm{Eu}$ sublayers, as shown by the black rectangle in Fig. 3(b). By comparing Figs. 2(f) and 2(g), we can see that the phonon Zeeman splitting generates a substantial topologically nontrivial phonon gap ($\approx 0.30$ THz) with a Chern number of $|C|=1$ featuring a pair of edge modes traversing the phonon gap, as shown in Fig. 3(h). This provides definitive evidence for intrinsic phonon Chern states.
@@
Discussion and conclusion. The intrinsic phonon Chern state, which hosts topologically protected edge phonon currents, holds significant potential for applications due to its operation without requiring strong external magnetic fields, Coriolis forces, or stress gradients [2, 49]. One potential advantage may be its ability to detect neutral particles by converting them into topological phonon edge currents. Neutral particles, especially dark matter particles, rarely participate in electromagnetic interactions, making their detection extremely challenging. However, if the energy transferred from the neutral particle to nucleus during the scattering or absorption falls within the first “conduction” band above the topological gap in the phonon dispersion, corresponding phonon modes can be excited [65, 66], driving the system out of equilibrium. As the system relaxes back to equilibrium, these high-frequency modes gradually shift to lower frequencies and necessarily cross the phonon Chern gap via topologically protected edge states, manifesting as a unidirectional edge phonon current immune to scattering [Fig. 4(a)]. This robust edge current is significantly more measurable than conventional delocalized bulk phonon excitations in materials. However, the operating temperature of detectors must be sufficiently low to prevent thermal excitation from overcoming the phonon gap, which would otherwise disrupt these edge currents. Thus, a large phonon Chern gap is highly desired to elevate the working temperature. Figure 4(b) lists phonon Chern gaps and transferred energy ranges of proposed candidate materials (see SM [44] for details), with transferred energy spanning approximately $20\sim 110\,\mathrm{meV}$. This energy range corresponds to low-energy dark matter which recently attracts much attention, therefore our work provides a potential detection mechanism for this category of dark matter candidates [65, 66].
Figure 4: Detection of neutral particles based on the intrinsic phonon Chern state. (a) Schematic illustrations of the detection process. Left: Neutral particles (red beam) excite phonon modes (yellow circles in the lower panel) in high-frequency bands. Right: These phonon modes gradually relax to lower-frequency regimes via topologically protected edge phonon channels (orange and green lines across the gap), generating robust unidirectional phonon edge currents (orange and green arrows). (b) Transferred energy ranges and phonon Chern gaps from proposed candidate materials [Tab. S2 in SM [44]].
In summary, we have established an ab initio theory enabling systematic predictions of EPC-induced magnetic properties of phonons and TRS-breaking phonon dispersions for insulators, semimetals and metals. Based on our theory, we construct the inertially decoupled model to seek intrinsic phonon Chern states in realistic materials, and propose candidates exhibiting EPC-induced large topological phonon gaps exceeding 0.3 THz. Furthermore, we demonstrate the potential capability of these topological phonon states to detect neutral particles through robust phonon edge currents. These advances could open broad pathways for investigating phonon magnetic effects in quantum materials.
@@ Acknowledgements.
The authors thank Lei Zhang and Zuowei Liu for valuable discussions. This work is partly supported by National Key Projects for Research and Development of China (Grants No. 2024YFA1409100, No. 2021YFA1400400 and No. 2022YFA1403602), the Natural Science Foundation of Jiangsu Province (No. BK20220032, BK20233001, BK20243011), the Natural Science Foundation of China (No. 12074181, No. 12104217, No. 12174182, and No. 92365203), and the e-Science Center of Collaborative Innovation Center of Advanced Microstructures.

### v4

@@ Abstract
Through ab initio calculations guided by the inertially decoupled model, we identified five hexagonal ($P6/mmm$) magnets exhibiting intrinsic phonon Chern states, namely, $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ [75], $\mathrm{EuGa_{2}}$, $\mathrm{GdGa_{2}}$ [76], and $\mathrm{CrB_{2}}$ [77, 78]. Remarkably, they all have large observable topological phonon gaps at the $\mathrm{K}$ point defined as $\Delta_{\text{ph}}=\Delta_{\text{z}}-\gamma$, where $\Delta_{\text{z}}=2|E_{\text{Z}}|$ is the phonon Zeeman splitting and $\gamma$ is the phonon linewidth resulting from EPC [79, 80], as summarized in Table I. Note that the Zeeman splitting should be larger than the phonon linewidth to produce an observable phonon gap.
@@
Here, we take $\mathrm{EuSi_{2}}$ as a representative example (see SM [58] for details on all candidate materials). The lattice structures of $\mathrm{EuSi_{2}}$ are illustrated in Figs. 4(a, b). Nonmagnetic $\mathrm{Si}$ atoms form a honeycomb-lattice layer, stacked alternately with a ferromagnetic $\mathrm{Eu}$ triangular-lattice layer, along the [001] direction. The atomic masses of $\mathrm{Si}$ and $\mathrm{Eu}$ are approximately 28u28\,u and 152u152\,u ($u$ is atomic mass unit), respectively, with each $\mathrm{Eu}$ atom carrying an out-of-plane magnetic moment of $7\,\mu_{B}$. As a result, $\mathrm{EuSi_{2}}$ can be aptly described by our proposed inertially decoupled model.
Figure 4: Lattice structure and TRS-breaking phonon dispersions of ferromagnetic $\mathrm{EuSi_{2}}$. (a,b) The front view (a) and the side view (b) of the lattice structure of $\mathrm{EuSi_{2}}$. (c) The Brillouin zones of bulk $\mathrm{EuSi_{2}}$ and the projected side plane. (d,e) The phonon dispersions of bulk $\mathrm{EuSi_{2}}$ without (d) and with (e) the phonon Zeeman splittings, respectively. The calculated PMMs at $K(K^{\prime})$ and $H(H^{\prime})$ point are $5.7\times 10^{-3}\mu_{B}$ and $2.3\times 10^{-2}\mu_{B}$, respectively. (f) The phonon dispersion of bulk $\mathrm{EuSi_{2}}$ under open-boundary conditions. (g,h) The phonon dispersion of the sandwiched slab enclosed by the black rectangle in (b) without (g) and with (h) the phonon Zeeman splittings, respectively. The PMM at $K(K^{\prime})$ point is $2.2\times 10^{-2}\mu_{B}$. (i) The phonon dispersion of the sandwiched slab under open-boundary conditions. The color gradient from red to blue represents the phonon angular momentum. Cyan and magenta colors in (f) and (i) label topological boundary modes residing on opposite boundaries.
@@
Our ab initio calculations reveal that EPC-enhanced PMM ($\sim 10^{-2}\mu_{B}$) induces phonon Zeeman splitting in the high-frequency phonon dispersion of ferromagnetic $\mathrm{EuSi_{2}}$ [see Figs. 4(d, e)], opening a sizable full topological gap ($\approx 0.35$ THz). Given that the Chern number remains $|C|=1$ for all 2D $k_{z}$ slices, $\mathrm{EuSi_{2}}$ realizes a 3D layered phonon Chern state. This is further supported by the presence of boundary modes in the phonon dispersions of $\mathrm{EuSi_{2}}$ under open-boundary conditions, as illustrated in Fig. 4(f). To achieve a 2D Chern state, we construct a dynamically-stable sandwiched slab featuring a $\mathrm{Si}$ sublayer between two $\mathrm{Eu}$ sublayers, as shown by the black rectangle in Fig. 4(b). By comparing Figs. 4(g) and 4(h), we can also see that the phonon Zeeman splitting generates a topologically nontrivial phonon gap ($\approx 0.30$ THz) with a Chern number of $|C|=1$ featuring a pair of edge modes traversing the phonon gap, as shown in Fig. 4(i). This provides definitive evidence for intrinsic phonon Chern states.
@@
Conclusion. Based on the linear response theory, we have established an ab initio theory, enabling systematic investigation of EPC-induced magnetic phonon spectra and phonon magnetic properties for a wide range of insulators, semimetals, and metals. This theoretical method is validated through the excellent agreement with the recent experiment results of magnetic Weyl semimetal $\mathrm{Co_{3}Sn_{2}S_{2}}$, providing a quantitative explanation of the phonon Zeeman splitting induced by magnetic order. We also propose an inertially decoupled model to seek intrinsic phonon Chern states in realistic materials, and predict $\mathrm{EuSi_{2}}$ family materials as candidates exhibiting large topological phonon gaps exceeding 0.3 THz, which could be applied to detect neutral particles, such as dark matter particles [58]. Therefore, our work paves a broad way for investigating phonon magnetic effects and phonon topology in quantum materials.
@@ Acknowledgements.
The authors thank Lei Zhang and Zuowei Liu for valuable discussions. This work is partly supported by National Key Projects for Research and Development of China (Grants No. 2024YFA1409100, No. 2021YFA1400400, No. 2022YFA1403602, and No. 2023YFA1407001), the Natural Science Foundation of Jiangsu Province (No. BK20253012, BK20252117, BK20233001, BK20243011, and BK20220032), the Natural Science Foundation of China (No. 12534007 and No. 92365203), and the e-Science Center of Collaborative Innovation Center of Advanced Microstructures.